(1) A large lightning bolt had a 20,500 A current and moved 32.5 C of charge. What was its duration (in ms)?

(2) A power transmission line is hung from metal towers with glass insulators having a resistance of 2.55 ✕ 109 Ω. What current (in mA) flows through the insulator if the voltage is 280 kV? (Some high-voltage lines are DC.)
Answer #1

1. Given i=20500 A ,charge =32.5

Time =c/i =32.5/20500= 1.58 milli second. =1.58×10^-3 seconds.

2.given r=2.55×10^9 ohm ,v =280 kv

From ohms law, v=Ir

I=280000/2.55×10^9 =1.09 milli Ampere.